REVTEX 4 Author’s Guide

American Physical Society^{†}

1 Research Road, Ridge, NY 11961

(Dated: May 2000)

Contents

1 Introduction 1 1.1 Submitting to APS Journals . . . . . . . . 1 1.2 Contact Information . . . . . . . . . . . . 1

2 Some L^{A}TEX 2ε Basics 2

2.1 Useful L^{A}TEX 2ε Markup . . . . . . . . . . 2

Fonts, User-defined macros, Symbols

2.2 Using L^{A}TEX 2ε packages with REVTEX . . 2

3 The Document Preamble 3 3.1 The documentclass line . . . . . . . . . . . 3 3.2 Loading other packages . . . . . . . . . . . 3

4 The Front Matter 3 4.1 Setting the title . . . . . . . . . . . . . . . 4 4.2 Specifying a date . . . . . . . . . . . . . . 4 4.3 Specifying authors and affiliations . . . . . 4

Footnotes for authors, Collaborations, Specifying first names and surnames

4.4 The abstract . . . . . . . . . . . . . . . . . 5 4.5 PACS codes . . . . . . . . . . . . . . . . . 5 4.6 Institutional report numbers . . . . . . . . 6 4.7 maketitle . . . . . . . . . . . . . . . . . . . 6

5 The body of the paper 6 5.1 Section headings . . . . . . . . . . . . . . 6 5.2 Paragraphs and General Text . . . . . . . 6 5.3 One-column vs. two-column . . . . . . . . 6 5.4 Cross-referencing . . . . . . . . . . . . . . 7 5.5 Acknowledgments . . . . . . . . . . . . . . 7 5.6 Appendices . . . . . . . . . . . . . . . . . 7

6 Math and Equations 7 6.1 Math in text . . . . . . . . . . . . . . . . . 7 6.2 Text in math . . . . . . . . . . . . . . . . 7 6.3 Displayed equations . . . . . . . . . . . . . 7 6.4 Numbering displayed equations . . . . . . 8 6.5 Cross-referencing displayed equations … 8

†Electronic address: revtex4@aps.org

6.6 Using the AMS packages amsfonts,

amssymb, and amsmath . . . . . . . . . . . 9 6.7 Bold symbols in math . . . . . . . . . . . 9

7 Footnotes 9

8 Citations and References 9 8.1 Author/Year (Non-numeric) Citations . . 10 8.2 Using BibTEX . . . . . . . . . . . . . . . . 10

9 Figures and Artwork 11

10Tables 11 10.1 Aligning on a decimal point . . . . . . . . 12 10.2 Footnotes in Tables . . . . . . . . . . . . . 12 10.3 Dealing with Long Tables . . . . . . . . . 12

11Placement of Figures, Tables, and Other Floats 12

12REVTEX 4 symbols and the revsymb package 13

13Other REVTEX 4 Features 13 13.1 Job-specific Override Files . . . . . . . . . 13

1. Introduction

This is the author’s guide to REVTEX 4, the preferred submission format for all APS journals. This guide is intended to be a concise introduction to REVTEX 4. The documentation has been separated out into smaller units to make it easier to locate essential information.

The following documentation is also part of the APS REVTEX 4 distribution. Updated versions of these will be maintained at the REVTEX 4 homepage located at

                                                http://publish.aps.org/revtex4/.

• REVTEX 4 Installation Guide

• APS Compuscript Guide

• REVTEX 4 Command and Options Summary

• REVTEX 4 BibTEXGuide • Differences between REVTEX 4 and REVTEX 3

• REVTEX 4 Frequently Asked Questions

This guide assumes a working REVTEX 4 installation. Please see the installation guide included with the distribution.

The REVTEX system for L^{A}TEX began its development in 1986 and has gone through three major revisions since then. All versions prior to REVTEX 4 were based on L^{A}TEX2.09 and, until now, REVTEX did not keep pace with the advances of the L^{A}TEX community and thus became inconvenient to work with. REVTEX 4 is designed to remedy this by incorporating the following design goals:

• Make REVTEX fully compatible with L^{A}TEX 2ε; it is now a L^{A}TEX 2ε document class, similar in function to the standard article class.

• Rely on standard L^{A}TEX 2ε packages for common tasks,

e.g, graphicx, color, and hyperref.

• Add or improve macros to support translation to tagged formats such as XML and SGML. This added markup will be key to enhancing the peer-review process and lowering production costs.

• Provide a closer approximation to the typesetting style used in Physical Review.

• Incorporate new features, such as hypertext, to make REVTEX a convenient and desirable e-print format.

• Relax the restrictions in REVTEX that had only been necessary for typesetting journal camera-ready copy.

To meet these goals, REVTEX 4 is a complete rewrite with an emphasis on maintainability so that it will be easier to provide enhancements.

The REVTEX 4 distribution includes both a template

(template.aps) and a sample document (apssamp.tex).

The template is a good starting point for a manuscript. In the following sections are instructions that should be sufficient for creating a paper using REVTEX 4.

1.1 Submitting to APS Journals

Authors using REVTEX 4 to prepare a manuscript for submission to Physical Review or Reviews of Modern Physics must also read the companion document APS Compuscript Guide distributed with REVTEX and follow the guidelines detailed there.

Further information about the compuscript program of the American Physical Society may be found at http:

//publish.aps.org/ESUB/.

1.2 Contact Information

Any bugs, problems or inconsistencies, should reported

to REVTEX support at mailto:revtex4@aps.org. Re-

ports should include information on the error and a small sample document that manifests the problem if possible (please don’t send large files!).

2. Some L^{A}TEX 2ε Basics

A primary design goal of REVTEX 4 was to make it as compatible with standard L^{A}TEX 2ε as possible so that authors may take advantage of all that L^{A}TEX 2ε offers. In keeping with this goal, much of the special formatting that was built in to earlier versions of REVTEX is now accomplished through standard L^{A}TEX 2ε macros or packages. The books in the bibliography provide extensive coverage of all topics pertaining to preparing documents under L^{A}TEX 2ε. They are highly recommended.

2.1 Useful L^{A}TEX 2ε Markup

L^{A}TEX 2ε markup is the preferred way to accomplish many basic tasks.

Fonts

Because REVTEX 4 is based upon L^{A}TEX 2ε, it inherits all of the macros used for controlling fonts. Of particular importance are the L^{A}TEX 2ε macros \textit, \textbf, \texttt for changing to an italic, bold, or typewriter font respectively. One should always use this macros rather than the lower-level TEX macros \it, \bf, and \tt. The L^{A}TEX 2ε macros offer improvements such as better italic correction and scaling in super- and subscripts for example. Table I summarizes the font selection commands in L^{A}TEX 2ε.

User-defined macros

L^{A}TEX 2ε provides several macros that enable users to easily create new macros for use in their manuscripts:

• \newcommand{⟨command⟩}[⟨narg⟩][⟨opt⟩]{⟨def ⟩}

• \newcommand*{⟨command⟩}[⟨narg⟩][⟨opt⟩]{⟨def ⟩}

• \renewcommand{⟨command⟩}[⟨narg⟩][⟨opt⟩]{⟨def ⟩}

• \renewcommand*{⟨command⟩}[⟨narg⟩][⟨opt⟩]{⟨def ⟩}

• \providecommand{⟨command⟩}[⟨narg⟩][⟨opt⟩]{⟨def ⟩}

• \providecommand*{⟨command⟩}[⟨narg⟩][⟨opt⟩]{⟨def ⟩} TABLE I: L^{A}TEX 2ε font commands

Text Fonts

Font command  Explanation
\textit{⟨text⟩}  Italics
\textbf{⟨text⟩}  Boldface
\texttt{⟨text⟩}  Typewriter
\textrm{⟨text⟩}  Roman
\textsl{⟨text⟩}  Slanted
\textsf{⟨text⟩}  Sans Serif
\textsc{⟨text⟩}  Small Caps
\textmd{⟨text⟩}  Medium Series
\textnormal{⟨text⟩} Normal Series
\textup{⟨text⟩}  Upright Series

Math Fonts

\mathit{⟨text⟩}  Math Italics
\mathbf{⟨text⟩}  Math Boldface
\mathtt{⟨text⟩}  Math Typewriter
\mathsf{⟨text⟩}  Math Sans Serif
\mathcal{⟨text⟩}  Calligraphic
\mathnormal{⟨text⟩} Math Normal
\bm{⟨text⟩}  Bold math for Greek letters

and other symbols

\mathfrak{⟨text⟩}^{a}  Fraktur
\mathbb{⟨text⟩}^{a}  Blackboard Bold

aRequires amsfonts class option

Here ⟨command⟩ is the name of the macro being defined, ⟨narg⟩ is the number of arguments the macro takes, ⟨opt⟩ are optional default values for the arguments, and ⟨def ⟩ is the actually macro definiton. \newcommand creates a new macro, \renewcommand redefines a previously defined macro, and \providecommand will define a macro only if it hasn’t been defined previously. The *-ed versions are an optimization that indicates that the macro arguments will always be “short” arguments. This is almost always the case, so the *-ed versions should be used whenver possible.

The use of these macros is preferred over using plain TEX’s low-level macros such as \def,\edef, and \gdef. APS authors must follow the APS Compuscript Guide when defining macros.

Symbols

L^{A}TEX 2ε has added some convenient commands for some special symbols and effects. These are summarized in Table II. See [3] for details.

2.2 Using L^{A}TEX 2ε packages with REVTEX

Many L^{A}TEX 2ε packages are available, for instance,

on CTAN at ftp://ctan.tug.org/tex-archive/
macros/latex/required/ and at ftp://ctan.tug.

TABLE II: L^{A}TEX 2ε commands for special symbols and effects

Command  Symbol/Effect

      \textemdash  —
      \textendash  –
      \textexclamdown  ¡
      \textquestiondown  ¿
      \textquotedblleft  “
      \textquotedblright  ”
      \textquoteleft  ‘
      \textquoteright  ’
      \textbullet  •
      \textperiodcentered  ·

      \textvisiblespace

      \textcompworkmark  Break a ligature

\textcircled{⟨char⟩}  Circle a character

org/tex-archive/macros/latex/contrib/ or may

be available on other distribution media, such as the

TEX Live CD-ROM http://www.tug.org/texlive.

Some of these packages are automatically loaded by REVTEX 4 when certain class options are invoked and are, thus, “required”. They will either be distributed with REVTEX or are already included with a standard L^{A}TEX 2ε distribution. The REVTEX 4 Installation Guide contains details about these packages.

Required packages are automatically loaded by REVTEX on an as-needed basis. Other packages should be loaded using the \usepackage command. To load the hyperref package, the document preamble might look like:

\documentclass{revtex}
\usepackage{hyperref}

Some common (and very useful) L^{A}TEX 2ε packages are a priori important enough that REVTEX 4 has been designed to be specifically compatible with them. A bug stemming from the use of one of these packages in conjunction with any of the APS journals may be reported by contacting REVTEX support.

AMS packages REVTEX 4 is compatible with and depends upon the AMS packages amsfonts, amssymb, and amsmath. In fact, REVTEX 4 requires use of these packages to accomplish some common tasks. See Section 6 for more.

array and dcolumn The array and dcolumn packages are part of L^{A}TEX’s required suite of packages. dcolumn is required to align table columns on decimal points (and it in turn depends upon the array package).

longtable longtable.sty may be used for large tables that will span more than one page. REVTEX 4 dynamically applies patches to longtable.sty so that it will work in two-column mode. hyperref hyperref.sty is a package by Sebastian

Rahtz that is used for putting hypertext links into L^{A}TEX 2ε documents. REVTEX 4 has hooks to allow e-mail addresses and URL’s to become hyperlinks if hyperref is loaded.

Other packages will conflict with REVTEX 4 and should be avoided. Usually such a conflict arises because the package adds enhancements that REVTEX 4 already includes. Here are some common packages that clash with REVTEX 4:

multicol multicol.sty is a package by Frank Mittelbach that adds support for multiple columns. In fact, early versions of REVTEX 4 used multicol.sty for precisely this. However, to improve the handling of floats, REVTEX 4 now has its own macros for two-column layout. Thus, it is not necessary to use

multicol.sty.

cite Donald Arseneau’s cite.sty is often used to provide support for sorting a \cite command’s arguments into numerical order and to collapse consecutive runs of reference numbers. REVTEX 4 has this functionality built-in already via the natbib package.

3. The Document Preamble

The preamble of a L^{A}TEX document is the set of commands that precede the \begin{document} line. It contains a \documentclass line to load the REVTEX 4 class (i. e., all of the REVTEX 4 macro definitions), \usepackage macros to load other macro packages, and other macro definitions.

3.1 The documentclass line

The basic formatting of the manuscript is controlled by setting class options using

\documentclass[⟨options⟩]{revtex4}. The macro
\documentclass replaces the \documentstyle macro of

L^{A}TEX2.09. The optional arguments that appear in the square brackets control the layout of the document. At this point, one only needs to choose a journal style (pra,

prb, prc, prd, pre, prl, prstab, and rmp) and either

preprint or twocolumn. Usually, one would want to use preprint for draft papers. twocolumn gives the Physical Review look and feel. A full list of class options is given in the REVTEX 4 Command and Options Summary.

3.2 Loading other packages

4. The Front Matter

After choosing the basic look and feel of the document by selecting the appropriate class options and loading in whatever other macros are needed, one is ready to move on to creating a new manuscript. After the preamble, be sure to put in a \begin{document} line (and put in an \end{document} as well). This section describes the macros REVTEX 4 provides for formatting the front matter of the article. The behavior and usage of these macros can be quite different from those provided in either REVTEX 3 or L^{A}TEX 2ε. See the included document Differences between REVTEX 4 and REVTEX 3 for an overview of these differences.

4.1 Setting the title

The title of the manuscript is simply specified by using the \title[⟨short title⟩]{title} macro. An optional short running title for display on the running heads of subsequent pages may be included as an optional argument.

4.2 Specifying a date

The \date{⟨date⟩} command outputs the date on the manuscript. Using \today will cause L^{A}TEX to insert the current date whenever the file is run:

\date{\today}

4.3 Specifying authors and affiliations

The macros for specifying authors and their affiliations have changed significantly for REVTEX 4. They have been improved to save labor for authors and in production. Authors and affiliations are arranged into groupings called, appropriately enough, author groups. Each author group is a set of authors who share the same set of affiliations. Author names are specified with the \author macro while affiliations (or addresses) are specified with the \affiliation macro. Author groups are specified by sequences of \author macros followed by \affiliation macros. An \affiliation macro applies to all previously specified \author macros which don’t already have an affiliation supplied.

For example, if Bugs Bunny and Roger Rabbit are both at Looney Tune Studios, while Mickey Mouse is at Disney World, the markup would be:

\author{Bugs Bunny}
\author{Roger Rabbit}
\affiliation{Looney Tune Studios}
\author{Mickey Mouse}
\affiliation{Disney World}

The default is to display this as

Bugs Bunny and Roger Rabbit Looney Tune Studios Mickey Mouse Disney World

This layout style for displaying authors and their affiliations is chosen by selecting the class option groupedaddress. This option is the default for all APS journal styles, so it does not need to be specified explicitly. The other major way of displaying this information is to use superscripts on the authors and affiliations. This can be accomplished by selecting the class option superscriptaddress. To achieve the display

Bugs Bunny,^{1} Roger Rabbit,^{1,2} and Mickey Mouse^{2}

1Looney Tune Studios 2Disney World

one would use the markup

\author{Bugs Bunny}
\affiliation{Looney Tune Studios}
\author{Roger Rabbit}
\affiliation{Looney Tune Studios}
\affiliation{Disney World}
\author{Mickey Mouse}
\affiliation{Disney World}

Note that REVTEX 4 takes care of any commas and and’s that join the author names together and font selection, as well as any superscript numbering. Only the author names and affiliations should be given within their respective macros.

There is a third class option, unsortedaddress, for controlling author/affiliation display. The default groupedaddress will actually sort authors into the approriate author groups if one chooses to specify an affiliation for each author. The markup:

\author{Bugs Bunny}
\affiliation{Looney Tune Studios}
\author{Mickey Mouse}
\affiliation{Disney World}
\author{Roger Rabbit}
\affiliation{Looney Tune Studios}

will result in the same display as for the first case given above even though Roger Rabbit is specified after Mickey Mouse. To avoid Roger Rabbit being moved into the same author group as Bugs Bunny, use the unsortedaddress option instead. In general, it is safest to list authors in the order they should appear and specify affiliations for multiple authors rather than one at a time. This will afford the most independence for choosing the display option. Finally, it should be mentioned that the affiliations for the superscriptaddress are presented and numbered in the order that they are encountered. These means that the order will usually follow the order of the authors. An alternative ordering can be forced by including a list of \affiliation commands before the first \author.

Footnotes for authors

Often one wants to specify additional information associated with an author such an e-mail address or an alternate affiliation. REVTEX 4 introduces several new macros just for this purpose. They are:

• \email[⟨optional text⟩]{e-mail address}

• \homepage[⟨optional text⟩]{URL}

• \altaffiliation[⟨optional text⟩]{affiliaton}

• \thanks{miscellaneous text}

In the first three, the optional text will be prepended before the actual information specified in the required argument. \email and \homepage each have a default text for their optional arguments (‘Electronic address:’ and ‘URL:’ respectively). The \thanks macro should only be used if one of the other three do not apply. Any author name can have multiple occurences of these four macros. Note that unlike the \affiliation macro, these macros only apply to the \author that directly precedes it. Any \affiliation must follow the other authorspecific macros. A typical usage might be as follows:

                                                  \author{Bugs Bunny}
                                                  \email[E-mail me at: ]{bugs@looney.com}
                                                  \homepage[Visit: ]{http://looney.com/}
                                                  \altaffiliation[Permanent address: ]
                                                                       {Warner Brothers}
                                                  \affiliation{Looney Tunes}

This would result in the footnote “E-mail me at:

                                                  bugs@looney.com, Visit: http://looney.com/, Perma-

nent address: Warner Brothers” being attached to Bugs Bunny. Note that:

• Only an e-mail address, URL, or affiliation should go in the required argument in the curly braces.

• The font is automatically taken care of.

• An explicit space is needed in the optional text if one is desired in the output.

• Use the optional arguments to provide customized text if there is a good reason to.

Each APS journal has a default behavior for the placement of footnotes on author names. The prb option puts all such footnotes at the start of the bibliography while the other journal styles display them on the first page. One can override a journal style’s default behavior by specifying explicitly the class option bibnotes (puts the footnotes at the start of the bibliography) or nobibnotes (puts them on the first page).

Collaborations

A collaboration name can be specified with the \collaboration macro. This is very similar to the \author macro, but it can only be used with the class op-

tion superscriptaddress. The \collaboration macro

should appear at the end of the list of authors. The collaboration name will be appear centered in parentheses between the list of authors and the list of affiliations. \collaboration can be followed by any of

\email, \homepage, or \thanks. Because collabora-

tions don’t normally have affiliations, one needs to follow the \collaboration and any associated \email,

\homepage, or \thanks macros with a \noaffiliation

macro. \noaffiliation may also be applied to an \author as well, but this is uncommon.

Specifying first names and surnames

Many APS authors have names in which either the surname appears first or in which the surname is made up of more than one name. To ensure that such names are accurately captured for indexing and other purposes, the \surname macro should be used to indicate which portion of a name is the surname. Similarly, there is a \firstname macro as well, although usage of \surname should be sufficient. If an author’s surname is a single name and written last, it is not necessary to use these macros. These macros do nothing but indicate how a name should be indexed. Here are some examples;

\author{Andrew \surname{Lloyd Weber}}
\author{\surname{Mao} Tse-Tung}

4.4 The abstract

An abstract for a paper is specified by using the abstract environment:

\begin{abstract}
Text of abstract
\end{abstract}

Note that in REVTEX 4 the abstract must be specified before the \maketitle command and there is no need to embed it in an explicit minipage environment. The current version of REVTEX 4 has the restriction that the abstract must fit on a single page.

4.5 PACS codes

APS authors are asked to supply suggested PACS codes with their submissions. The \pacs macro is provided as a way to do this:

                                                  \pacs{23.23.+x, 56.65.Dy}

The actual display of the PACS numbers below the abstract is controlled by two class options: showpacs and noshowpacs. In particular, this is now independent of the preprint option.

4.6 Institutional report numbers

Institutional report numbers can be specified using the \preprint macro. These will be displayed in the upper lefthand corner of the first page. Multiple \preprint macros maybe supplied.

4.7 maketitle

After specifying the title, authors, affiliations, abstract, PACS codes, and report numbers, the final step for formatting the front matter of the manuscript is to execute the \maketitle macro by simply including it:

                                                  \maketitle

The \maketitle macro must follow all of the macros listed above. The macro will format the front matter in accordance with the various class options that were specified in the \documentclass line (either implicitly through defaults or explicitly).

5. The body of the paper

For typesetting the body of a paper, REVTEX 4 relies heavily on standard L^{A}TEX 2ε and other packages (particulary those that are part of AMSTEX). Users unfamiliar with these packages should read the following sections carefully.

5.1 Section headings

Section headings are input as in L^{A}TEX. The output is similar, with a few extra features.

Four levels of headings are available in REVTEX:

\section[⟨short title⟩]{⟨title text⟩}

                                                  \subsection{⟨title text⟩}
                                                  \subsubsection{⟨title text⟩}
                                                  \paragraph{⟨title text⟩}

Provide the ⟨short title⟩ if needed for the sake of the running header (required only by some journal substyles). Use the starred form of the command to suppress the

automatic numbering; e.g.,

                                                  \section*{Introduction}

To label a section heading for cross referencing use the \label{⟨key⟩} command after the heading; e.g.,

\section{Introduction}
\label{sec:intro}

In the some journal substyles, such as those of the APS, all text in the \section command is automatically set uppercase. If a lowercase letter is needed, use \lowercase{x}. For example, to use “He” for helium in a \section{⟨title text⟩} command, type H\lowercase{e} in {⟨title text⟩}.

Use \protect\\ to force a line break in a section heading. (Fragile commands must be protected in section headings and captions, and \\ is a fragile command.)

5.2 Paragraphs and General Text

Paragraphs always end with a blank input line. Because TEX automatically calculates linebreaks and word hyphenation in a paragraph, it is not necessary to force linebreaks or hyphenation. Of course, compound words may should still be explicitly hyphenated, e.g., “authorprepared copy.”

Use directional quotes for quotation marks around quoted text (``xxx''), not straight double quotes ("xxx"). For opening quotes, use one or two backquotes; for closing quotes, use one or two forward quotes (apostrophes).

5.3 One-column vs. two-column

One of the hallmarks of Physical Review is its twocolumn formatting and so one of the REVTEX 4 design goals is to make it easier to acheive the Physical Review look and feel. In particular, the twocolumn option will take care of formatting the front matter (including the abstract) as a single column. REVTEX 4 has its own built-in two-column formatting macros to provide wellbalanced columns as well as reasonable control over the placement of floats in either one- or two-column modes.

Occasionally it is necessary to change the formatting from two-column to one-column to better accomodate very long equations that are more easily read when typeset to the full width of the page. This is accomplished using the widetext environment:

\begin{widetext}
long equation goes here
\end{widetext}

In two-column mode, this will temporarily return to one-column mode, balancing the text before the environment into two short columns, and returning to twocolumn mode after the environment has finished. The widetext environment has no effect on the output under the preprint class option because this already uses one-column formatting.

Use of the widetext environment should be restricted to the bare minimum of text that needs to be typeset this way. However short pieces of paragraph text and/or math between nearly contiguous wide equations should be incorporated into the surrounding wide sections. REVTEX 4 works quite hard to get good balancing behavior, but can run into trouble if the widetext occurs too close to a page boundary. Symptoms of this problem include blocks of text disappearing or being moved to a strange location.

5.4 Cross-referencing

REVTEX inherits the L^{A}TEX 2ε features for labeling and cross-referencing section headings, equations, tables, and figures. This section contains a simplified explanation of these cross-referencing features. The proper usage in the context of section headings, equations, tables, and figures is discussed in the appropriate sections.

Cross-referencing depends upon the use of “tags,” which are defined by the user. The \label{⟨key⟩} command is used to identify tags for REVTEX. Tags are strings of characters that serve to label section headings, equations, tables, and figures that replace explicit, byhand numbering.

Files that use cross-referencing (and almost all manuscripts do) need to be processed through REVTEX at least twice to ensure that the tags have been properly linked to appropriate numbers. If any tags are added in subsequent editing sessions, L^{A}TEX will display a warning message in the log file that ends with ...

Rerun to get cross-references right. Running the

file through REVTEX again (possibly more than once) will resolve the cross-references. If the error message persists, check the labels; the same {⟨key⟩} may have been used to label more than one object.

Another L^{A}TEX warning is There were undefined references, which indicates the use of a key in a \ref without ever using it in a \label statement.

REVTEX performs autonumbering exactly as in standard L^{A}TEX. When the file is processed for the first time, L^{A}TEX creates an auxiliary file (with the .aux extension) that records the value of each ⟨key⟩. Each subsequent run retrieves the proper number from the auxiliary file and updates the auxiliary file. At the end of each run, any change in the value of a ⟨key⟩ produces a L^{A}TEX warning message.

5.5 Acknowledgments

Use the acknowledgments environment for an acknowledgments section. Depending on the journal substyle, this element may be formatted as an unnumbered section title Acknowledgments or simply as a paragraph. Please note the spelling of “acknowledgments”.

\begin{acknowlegments}
The authors would like to thank...
\end{acknowlegments}

5.6 Appendices

The \appendix command signals that all following sections are appendices, so \section{⟨title text⟩} after \appendix will set {⟨title text⟩} as an appendix heading (an empty {⟨title text⟩} is permitted). For a single appendix, use a \section*{⟨title text⟩} command to suppress the appendix letter in the section heading.

6. Math and Equations

6.1 Math in text

Not surprisingly, REVTEX uses the TEX math $ delimiters for math embedded in text. For example, $a^{z}$ give a^{z}. Within math mode, use ^{⟨math⟩} for superscripts and _{⟨math⟩} for subscripts. If the braces after the ^are omitted, TEX will superscript the next token (generally a single character or command). Thus it is safest to use explicit braces {}.

As with text, math should not require extensive explicit vertical or horzontal motion commands, because TEX calculates math spacing itself automatically. In particular, explicit spacing around relations (e.g., =) or operators (e.g., +) should be unnecessary. These suggestions notwithstanding, some fine-tuning of math is required in specific cases, see Chapter 18 in the TEXbook[1].

6.2 Text in math

There are times when normal, non-italic text needs to be inserted into a math expression. The \text{⟨text⟩} command is the preferred method of accomplishing this. It produces regular text and scales correctly

in superscripts: $y=x \text{ for } x_{\text{e-p}}$

gives “y = x for x_{e-p}”. To use the \text command, the amsmath package must be loaded: include a \usepackage{amsmath} command in the document preamble.

Other common alternatives may be less desirable. Using the standard L^{A}TEX 2ε \mbox{⟨text⟩} will give normal text, including a hyphen, but will not scale correctly in superscripts: $x_{\mbox{e-p}}$ gives “xe-p”. The \rm command only switches to Roman font for math letters. It does not, for example, handle hyphens correctly: $$x_{\rm{e-p}}$ gives “x_{e−p}”. But note that \textrm, it does work: $x_{\textrm{e-p}}$ gives “x_{e-p}”.

6.3 Displayed equations

Equations are set centered in the column width or flush left depending on the selected journal substyle.

For the simplest type of displayed equation, a numbered, one-line equation, use the equation environment. REVTEX takes care of the equation number—the number will be set below the equation if necessary. Use \[… \] for a single, one-line unnumbered display equation.

Use the eqnarray environment when more than one consecutive equation occurs, putting each equation in a separate row of the environment, and using \nonumber before the row end (\\) to suppress the equation number where necessary. If the equations are related to each other, align each on the respective relation operator (such as =).

When an equation is broken over lines or is continued over multiple relation operators, it is called a multi-line or continued equation, respectively; here, too, use the eqnarray environment.

For a continued equation, align each row on the relation operator just as with multiple equations, and use the \nonumber command to suppress auto-numbering on broken lines. Also, use the starred form of the row end (\\*) to prevent a pagebreak at that juncture.

Short displayed equations that can appear together on a single line separated by \qquad space may be placed in a single equation environment.

As explained in Section 5.3, occasionally in two-column mode a long equation, in order to fit it in the narrow column width, would need to be broken into so many lines that it would affect readibility. Set it in a wide column using the widetext environment. Then return to the normal text width as soon as possible.

The sample file apssamp.tex illustrates how to obtain each of the above effects.

6.4 Numbering displayed equations

REVTEX 4 automatically numbers equations. For single-line and multi-line equations, use the equation and eqnarray environments as described above. For unnumbered single-line equations, use the \[… \] construction. The command \nonumber will suppress the numbering on a single line of an eqnarray. For a multiline equation with no equation numbers at all, use the eqnarray* environment.

A series of equations can be a labeled with a lettered sequence, e.g., (3a), (3b), and (3c), by putting the respective equation or eqnarray environment within a subequations environment. The amsmath package is required for this.

Use the command \tag{⟨number⟩} to produce an idiosyncratic equation number: (1′), for example. Numbers assigned by \tag are completely independent of REVTEX’s automatic numbering. The package amsmath is required if for using the \tag command.

To have REVTEX reset the equation numbers at the start of each section, use the eqsecnum class option in the document preamble.

See the sample file apssamp.tex for some examples.

6.5 Cross-referencing displayed equations

To refer to a numbered equation, use the \label{⟨key⟩} and \ref{⟨key⟩} commands. The \label{⟨key⟩} command is used within the referenced equation (on the desired line of the eqnarray, if a multi-line equation):

\begin{equation}
A=B \label{pauli}
\end{equation}
... It follows from Eq.~(\ref{pauli})
that this is the case ...
\begin{eqnarray}
A & = &B,\label{pauli2}\\
A'& = &B'
\end{eqnarray}

gives

A = B  (1)

... It follows from Eq. (1) that this is the case ...

A = B,  (2)
A′ = B′  (3)

Please note the parentheses surrounding the \ref command. These are not provided automatically and, thus, must be explicitly incorporated.

Numbers produced with \tag can also be crossreferenced by adding a \label command after the \tag command.

Using a \label after \begin{subequations} to ref-

erence the general number of the equations in the subequations environment. For example, if

\begin{subequations}
\label{allequations} % notice location
\begin{eqnarray}
E&=&mc^2,\label{equationa}
\\
E&=&mc^2,\label{equationb}
\\
E&=&mc^2,\label{equationc}
\end{eqnarray}
\end{subequations}

gives the output

E = mc^{2},  (4a)
E = mc^{2},  (4b)
E = mc^{2},  (4c)

then Eq.~(\ref{allequations}) gives “Eq. (4)”.

Note: incorrect cross-referencing will result if \label is used in an unnumbered single-line equation (i.e., within the \[ and \] commands), or if \label is used on a line of an eqnarray that is not being numbered (i.e., a line that has a \nonumber).

6.6 Using the AMS packages amsfonts, amssymb, and

                       amsmath

The American Mathematical Society’s AMSL^{A}TEX packages provided extra fonts, symbols, and math markup that are quite convenient. REVTEX 4 supports the use of these packages directly. To use the amsfonts, amssymb, and amsmath class options, AMSL^{A}TEX (and perhaps the additional AMS fonts) will need to be installed (see the REVTEX 4 Installation Guide).

There are two class options for accessing the AMS

fonts: amsfonts and amssymb. The amsfonts option de-

fines the \mathfrak and \mathbb commands to switch to the Fraktur and Blackboard Bold fonts, respectively. These fonts are selected with the \mathfrak and \mathbb font-switching commands: ${\mathfrak{G}}$ gives a Fraktur “𝔊” and ${\mathbb{Z}}$ gives a Blackboard Bold “ℤ”. REVTEX does not currently support the use of the extra Euler fonts (the AMS fonts starting with eur or eus) or the Cyrillic fonts (the AMS fonts starting with w).

The amssymb class option gives all the font capabilities of the amsfonts class option and further defines the commands to get the symbols shown in the REVTEX 4 Command and Options Summary, which contains examples of the symbols and for instructions on use. These characters will scale correctly in superscripts and heads.

6.7 Bold symbols in math

REVTEX 4 uses the standard L^{A}TEX 2ε Bold Math (bm) package as the basis for creating bold symbols in math mode. As usual, this requires an explicit \usepackage{bm} in the document preamble. The command \bm{⟨symbol⟩} makes {⟨symbol⟩} bold in math mode, ensuring that it is the correct size, even in superscripts. If the correct font in the correct size is not available then result is the {⟨symbol⟩} set at the correct size in lightface and a L^{A}TEX 2ε warning that says “No boldmath typeface in this size… ”. Most bold special characters will require that the AMS fonts be installed and the amsfonts class option be invoked.

\bm is the proper means to get bold greek characters— upper- and lowercase—and other symbols. The following will come out bold with \bm: normal math italic letters, numbers, Greek letters (uppercase and lowercase), small bracketing and operators, and \mathcal. Fraktur characters will come out bold in a \bm; however, Blackboard Bold requires using the \mathbbcommand rather than \bm. The amsfonts option adds support for bold math letters and symbols in smaller sizes and in superscripts when a \bm{⟨symbol⟩} is used. For example, $\pi^{\bm{\pi}}$ gives a bold lowercase pi in the superscript position: ^{ππ}.

Note that \bm{⟨math⟩} is a fragile command and, thus, should be preceded by \protect in commands with moving arguments.

7. Footnotes

L^{A}TEX’s standard \footnote command is available in REVTEX 4. The footnote text can either appear at the bottom of a page or as part of the bibliography (in order, after the rest of the references). Specific journal options will select one of these two methods. Note that in the latter case, the argument of the \footnote command is a moving argument in the sense of the L^{A}TEX User’s Guide & Reference Manual, Appendix C.1.3: any fragile command within that argument must be preceded by a

\protect command.

The \footnote macro should not be used in the front matter for indicating author/affiliation relationships or to provide additional information about authors (such as an e-mail address). See Section 4.3 for the proper way to do this.

Finally, footnotes that appear in tables behave differently. They will be typeset as part of the table itself. See Section 10.2 for details.

8. Citations and References

In keeping with the REVTEX 4 design goal of making it easier to extract tagged information from a manuscript, new macros and BibTEX style files have been added to provide better markup. Furthermore, these have been built upon some widely-used L^{A}TEX packages in line with the design goal of making use of the existing packages where possible. The two main external packages that are of concern here are Patrick Daly’s natbib citation package and his custombib tool kit for building new BibTEX style files.

From an author’s point of view, all this means is that a proper REVTEX 4 installation requires having natbib (version 7 or higher) installed. It also means that the full set of natbib functionality is available from within REVTEX 4 (but see the APS Compuscript Guide for restrictions).

As in standard L^{A}TEX, references are cited in text using the \cite{⟨key⟩} command and are listed in the bibliography using the \bibitem{⟨key⟩} command. The \cite macro enables REVTEX 4 to automatically number the references in the manuscript. A typical example might be:

                                                  String theory\cite{GSW} attempts to provide a
                                                  theory of everything.

The corresponding \bibitem would be:

                                                  \bibitem{GSW} M.~Greene, J.~Schwarz, and
                                                  E.~Witten, \textit{Superstring Theory:
                                                  Introduction}, (Cambridge University
                                                  Press, London, 1985).

Journals differ in how the \cite will be displayed. Most APS journals display the citation in-line, as a number, enclosed in square brackets, e. g., “String theory[1] attempts. …” Other journals (most notably Physical Review B) instead use a number in a superscript: “String theory^{1} attempts. …” Selecting the journal substyle using a class option (such as prb) will invoke the appropriate style. In journal substyles using superscripts, the macro the \onlinecite{⟨key⟩} is necessary to get the number to appear on the baseline. For example, “String theory (see, for example, Ref.~\onlinecite{GSW})” will give the output “String theory (see, for example, Ref. 1).” A \cite command with multiple keys is formatted with consecutive reference numbers collapsed; e.g., [1,2,3,5] will be output as [1–3,5]. To split the list over more than one line, use a % character immediately following a comma:

                                                  . . .  \cite{a,b,c,d,e,f,%
                                                  g,h,i,j,k,l,m,n,o,p,q,r,s,t,u,v,w,x,y,z}

The % avoids unwanted spaces.

8.1 Author/Year (Non-numeric) Citations

Reviews of Modern Physics uses a citation style based on the first author’s last name and the year of the reference rather than a simple number. Support for this style of citing references is the primary reason REVTEX 4 uses the natbib package. natbib uses an optional argument to the \bibitem macro to specify what text to use for the \cite text:

\bibitem[⟨short-name⟩(⟨year⟩)⟨long-name⟩]

where ⟨short-name⟩ is the author name used in a parenthetical citation, ⟨long-name⟩ that used in a textual citation, and ⟨year⟩ is the year. More concretely, the \bibitem example above would appear as

                                                  \bibitem[Greene et al.(1985)Green, Schwarz, and
                                                  Witten]{GSW} M.~Greene, J.~Schwarz, and

E.~Witten, \textit{Superstring Theory:
Introduction}, (Cambridge University Press,
London, 1985).

When the citation constitutes part of the grammar of the sentence, the \textcite{⟨key⟩} command may be used (analogous to the \onlinecite command above). Both \textcite and \onlinecite are built upon natbib’s rich repertoire of macros (\citep, \citet, etc.). These macros are available in REVTEX 4; however, APS authors must follow the REVTEX 4 Compuscript Guide guidelines regarding natbib’s macros.

8.2 Using BibTEX

The \bibitem entries can be coded by hand as above, of course, but the use of BibTEX with the new style files provided with REVTEX 4 makes it particularly simple to generate marked-up references that can, for instance, take advantage of packages like hyperref for linking. They also save the trouble of having to specify formatting like the italics for the book title in the above example. And, for those wishing to use author/year citations, BibTEXwill automatically generate the appropriate optional arguments for the \bibitem commands.

BibTEX is an adjunct to L^{A}TEX that aids in the preparation of bibliographies. BibTEX allows authors to build up a database or collection of bibliography entries that may be used for many manuscripts. A BibTEX style file then specifies how to transform the entries into a proper \bibitem for a particular journal. Here we give a brief summary of how to get started with BibTEX. More details can be found in the REVTEX 4 BibTEX Guide included in the distribution or in the LaTeX books listed in the references.

Selecting a journal style by using an appropriate class option will automatically select the correct BibTEX style file. For all of the Physical Review journals, this is apsrev.bst, while for Reviews of Modern Physics this is apsrmp.bst. The selection can be overridden by specifying an alternative .bst file using the standard L^{A}TEX 2ε\bibliographystyle macro. This must appear in the preamble before the \begin{document} line in REVTEX 4 (this differs from standard L^{A}TEX).

The BibTEX database files will contain entries such as:

@Book{GSW,
author=``M. Greene, J. Schwarz, E. Witten'',
title=``Superstring theory: Introduction'',
publisher=``Cambridge University Press'',
address=``London'',
year=``1985''
}

There are entry formats for articles, technical reports, eprints, theses, books, proceedings, and articles that appear in books or proceedings. The apsrev and apsrmp styles provided with REVTEX 4 also allows URL’s and e-print identifiers to be specified for any of the different entry types. There is also add a “collaboration” field that can be used in addition to “author”.

To actually create the bibliography in the manuscript, the \bibliography{⟨bib files⟩} macro is used. Here ⟨bib files⟩ is a comma-separated list of BibTEX bibliography database files, each with the .bib extension. The \bibliography macro should be placed at the location where the references should appear (usually after the main body of the paper). When the manuscript is processed with L^{A}TEX for the first time, the keys corresponding for the \cite macros used in the manuscript are written out to the .aux file. Then BibTEX should be run (if the manuscript is called man.tex, the command would be “bibtex man.” This will produce a .bbl file containing all of the \bibitem’s for the manuscript. Subsequent runs of L^{A}TEX 2ε will call this file in to resolve the references. L^{A}TEX 2ε should be run repeatedly until all references are resolved.

The BibTEX-produced \bibitem’s created using the apsrev and apsrmp appear considerably more complex than the example given above. This is because the style

files add in \bibinfo, \bibnamefont, \eprint, and \url

macros for specifying additional formatting and tagging. The \bibinfo macro is mostly a do-nothing macro that serves merely to tag the information with the field information from the original entry in the BibTEX database. The \eprint and \url macros can be used to create the appropriate hyperlinks in target formats such as PDF.

For more information on using BibTEX with L^{A}TEX, see Sections 4.3.1 and C.11.3 of the L^{A}TEX User’s Guide & Reference Manual[2], Section 13.2 of [4], or the online

BibTEX manual http://ctan.tug.org/tex-archive/
biblio/bibtex/distribs/doc/btxdoc.tex.

9. Figures and Artwork

Figures may be included into a REVTEX 4 manuscript by using the standard L^{A}TEX 2ε macros. It should be noted that L^{A}TEX 2ε includes several powerful packages for including the files in various formats. The two main packages are graphics and graphicx. Both offer a macro called \includegraphics[⟨args⟩]{⟨filename⟩}; they mainly differ in how arguments for controlling figure placement (e. g., scaling and rotating) are passed to the \includegraphics. Old L^{A}TEX2.09 interfaces such as epsfig have been re-implemented on top of these packages. For more information on the enhancements of the graphicx package, see [5]

or ftp://ctan.tug.org/tex-archive/macros/latex/
required/graphics/grfguide.ps. REVTEX 4 no

longer has the epsf class option, though the epsfig package provides a similar interface.

The figure environment should be used to add a caption to the figure and to allow L^{A}TEX to number and place the figures where they fit best. L^{A}TEX will label and automatically number the captions FIG. 1, FIG. 2, etc. For example:

\begin{figure}
 \includegraphics{fig1.eps}
 \caption{Text of first caption.}
 \label{fig1}
\end{figure}

Note how the \label{⟨key⟩} command is used to crossreference figures in text. The \label{⟨key⟩} command should be inserted inside or after the figure caption, before the end of the figure environment. As usual, the \ref{⟨key⟩} macro can then by used to refer to the label: “As depicted in FIG.~\ref{fig1}… ”.

10. Tables

Tables are very similar to figures. They should be input using the table environment as detailed below, and L^{A}TEX will label and number the captions TABLE 1, TA-BLE 2, etc. (or in whatever format required by the chosen journal substyle). Tables without captions won’t be numbered.

Each table must begin with \begin{table}, end with \end{table}. A a caption can be specified using the \caption{⟨text⟩} command. To refer to the table via cross-referencing, a \label{⟨key⟩} command should follow the \caption. Use the \ref{⟨key⟩} command to cite tables in text. The table environment will set the table to the width of the column. Thus, in two-column mode, the table will confined to a single column. To set a table to the full width of the page, rather than the column, use the table* environment.

The heart of the table is the tabular environment. This will behave for the most part as in standard L^{A}TEX 2ε (please refer to Section 3.6.3 and Appendix C.10.2 of the L^{A}TEX User’s Guide & Reference Manual for more details about the tabular environment). Note that REVTEX 4 no longer automatically adds double (Scotch) rules around tables. Nor does the tabular environment set various table parameters for column spacing as before. Instead, a new environment ruledtabular provides this functionality. This environment should surround the tabular environment:

\begin{table}
\caption{....}
\label{tab:....}
\begin{ruledtabular}
\begin{tabular}
...
\end{tabular}

                                                  \end{ruledtabular}
                                                  \end{table}

A basic table looks as follows:

                                                  \begin{table}
                                                  \caption{Text of table caption.}
                                                  \label{tab:example}
                                                  \begin{ruledtabular}
                                                  \begin{tabular}{ll}
                                                    Heading 1 & Heading 2\\
                                                    Cell 1 & Cell 2\\
                                                  \end{tabular}
                                                  \end{ruledtabular}
                                                  \end{table}

The quasitable environment is no longer in REVTEX 4. The standard tabular environment can be used instead because it no longer puts in the double rules.

10.1 Aligning on a decimal point

Numerical columns should align on the decimal point (or decimal points if more than one is is present). This is accomplished by again using a standard L^{A}TEX 2ε package, dcolumn which must be loaded in the manuscript’s preamble:

                                                  \usepackage{dcolumn}

Once this package is loaded, the column specifier ‘d’ can be used in the table’s tabular{⟨preamble⟩} enviroment preamble. The ‘d’ should be used for simple numeric data with a single decimal point. The entry of a d column is typeset in math mode; do not insert any $ math delimiters into a ‘d’ column. Items without a decimal point are simply set in math mode, centered. If text is required in the column, use \text or \mbox as appropriate. If multiple decimal points are present then the last is used for alignment. To escape from the ‘d’ column use

                                                  \multicolumn as usual. See the sample file apssamp.tex

for examples.

10.2 Footnotes in Tables

Footnotes in a table are labeled a, b, c, etc. They can be specified by using the L^{A}TEX \footnote command.

                                                  Furthermore, \footnotemark and \footnotetext can

be used so that multiple entries can to refer to the same footnote. The footnotes for a table are typeset at the bottom of the table, rather than at the bottom of the page or at the end of the references. The arguments for

                                                  \footnotemark and \footnotetext should be numbers

1, 2, …. The journal style will convert these to letters. See sample file apssamp.tex for examples and explanations of use.

10.3 Dealing with Long Tables

By default, tables are set in a smaller size than the text body (\small). The \squeezetable declaration makes the table font smaller still (\scriptsize). Thus, putting the \squeezetable command before the \begin{tabular} line in a table will reduce the font size. If this isn’t sufficient to fit the table on a page, the standard L^{A}TEX 2ε longtable package may be used.

Under REVTEX 3, tables automatically break across pages; REVTEX 4 provides some of this functionality. However, this requires adding the table a float placement option of [h] (meaning put the table “here”) to the \begin{table} command. Furthermore, if the ruledtabular environment is not used, the \begin{tabular} command must be given an optional [v] argument. Tables set in the table* environment ignore any float placement options.

Long tables are more robustly handled by using the longtable.sty package included with the standard L^{A}TEX 2ε distribution (put \usepackage{longtable} in the preamble). This package gives precise control over the layout of the table. The REVTEX 4 package contains patches that enable the longtable package to work in two-column mode. Of course, a table set in two-column mode needs to be narrow enough to fit within the column. Otherwise, the columns may overlap. A longtable that is too wide for the narrow columns can be set within a widetext environment. (Note that a standard table should never be enclosed in a widetext environment - use the table* enviroment instead.) For more documentation on the longtable environment and on the package options of the longtable package, please see the doc-

umentation available at ftp://ctan.tug.org/macros/
latex/required/tools/longtable.dtx or refer to [4].

Alternatively, as in REVTEX 3.1, there is support in REVTEX 4 for automatic table breaking. REVTEX 4 will break a tabular enviroment across page breaks. This can only happen if the enclosing environment (table or table*) has either the ‘h’ float placement directive specified and the tabular environment either has an optional argument [v] specified or it is enclosed within a

ruledtabular environment.

11. Placement of Figures, Tables, and Other Floats

By default, figures and tables (and any other “floating” environments defined by other packages) float to the top or bottom of the page using the standard L^{A}TEX float placement mechanism. Initially, each figure or table environment should be put immediately following its first reference in the text; this will usually result in satisfactory placement on the page. An optional argument for either environment adjusts the float placement. For example:

\begin{figure}[⟨placement⟩]

…

\end{figure}

where ⟨placement⟩ can be any combination of htbp!, signifying “here”, “top”, “bottom”, “page”, and “as soon as possible”, respectively. The same placement argument may be added to a \begin{table}. For more details about float placement, see the instructions in the L^{A}TEX User’s Guide & Reference Manual, Appendix C.9.1.

In two-column mode, a page may contain both a widetext environment and a float. The placement of floats in such a situation can be very tricky and REVTEX 4 may not always be able to automatically put the float in the optimal place. For instance, a float may be placed at the bottom of a column just before the widetext begins. To workaround this, try moving the float environment below the widetext environment. Trying alternative ⟨placements⟩ may also alleviate the problem.

Figures and tables are normally set to the width of the column in which they are placed. This means that in two-column mode, the float will be placed in a single, narrow column. For wide figures or tables, the \figure* or \table* environments should be used instead. This will place the float across both columns (the float will appear either at the top or the bottom of the page in this case). The widetext environment should not be used to accomplish this effect.

REVTEX 4 offers an additional possibility for placing the floats. By using the either the endfloats or the endfloats* class option all floats may be held back (using an external file) and then set elsewhere in the document using the the commands \printtables and \printfigures, placed where the tables and figures are to be printed (usually at the end of the document). (This is similar to the standard \printindex command). Using a *-form of the commands (\printfigures* and \printtables*) will begin the figures or tables on a new page. Alternatively, the option endfloats* may be used to change the behavior of the non-*-forms so that every float will appear on a separate page at the end.

Without one of the endfloats class options, these float placement commands are silently ignored, so it is always safe to use them. If one of the endfloats class options is given, but the \printtables command is missing, the tables will be printed at the end of the document. Likewise, if \printfigures is missing, the figures will be printed at the end of the document. Therefore it is also safe to omit these commands as long as REVTEX’s default choices for ordering figures and tables are satisfactory.

The endfloats option (or perhaps some journal substyle that invokes it), requires explicit \begin{figure},

\end{figure}, \begin{table}, and \end{table} lines.

In particular, do not define typing shortcuts for table and figure environments, such as TABLE III: Special REVTEX 4 symbols, accents, and boldfaced parentheses defined in revsymb.sty

\lambdabar         λ  \openone  11
\altsuccsim  ≻_{∼}  \altprecsim  ≺_{∼}
\alt  ≲  \agt  ≳
\tensor x  x^{↔}  \overstar x  x^{∗}
\loarrow x  x^{←}  \roarrow x  x^{→}
\biglb ( \bigrb)  ((())) \Biglb ( \Bigrb)  ( ( () ) )

\bigglb ( \biggrb) ( ( () ) ) \Bigglb ( \Biggrb ) ( ( () ) )

\def\bt{\begin{table}}% Incompatible!
\def\et{\end{table}}%

12. REVTEX 4 symbols and the revsymb package

Symbols made available in earlier versions of REVTEX are defined in a separate package, revsymb, so that they may be used with other classes. This might be useful if, say, copying text from a REVTEX document to a non-REVTEX document. REVTEX 4 automatically includes these symbols so it is not necessary to explicitly call them in with a \usepackage statement.

Table III summarizes the symbols defined in this package. Note that \overcirc, \overdots, and \corresponds are no longer in REVTEX 4. Use \mathring (standard in L^{A}TEX 2ε), \dddot (with the amsmath package loaded), and \triangleq (with the amssymb class option) respectively.

13. Other REVTEX 4 Features

13.1 Job-specific Override Files

REVTEX 4 allows manuscript-specific macro definitions to be put in a file separate from the main TEX file. One merely creates a file with the same basename as the TEX file, but with the extension ‘.rty’. Thus, if the TEX file is names man.tex, the macro definitions would go in man.rty. Note that the .rty file should be in the same directory as the TEX file. APS authors should follow the guidelines in the REVTEX 4 Compuscript Guide when submitting.

[1] D.E. Knuth, The TEXbook, (Addison-Wesley, Reading, MA, 1986).

[2] L. Lamport, L^{A}TEX, a Document Preparation System, (Addison-Wesley, Reading, MA, 1996).

[3] H. Kopka and P. Daly, A Guide to L^{A}TEX 2ε, (Addison-Wesley, Reading, MA, 1995).

[4] M. Goosens, F. Mittelbach, and A. Samarin, The L^{A}TEX Companion, (Addison-Wesley, Reading, MA, 1994).

[5] M. Goossens, S. Rahtz, and F. Mittelbach, The L^{A}TEX Graphics Companion, (Addison-Wesley, Reading, MA, 1997).

[6] S. Rahtz, M. Goossens, et al.,The L^{A}TEX Web Companion, (Addison-Wesley, Reading, MA, 1999).